Find the first five outputs of the following recursively defined function:
f(1)=8,(thats given) f(n+1)=-f(n)+6
its not really a permutation but i didnt know what category to put it in

Answer by stanbon(75887)

Find the first five outputs of the following recursively defined function:
f(1)=8,(thats given) f(n+1)=-f(n)+6
---------------
f(1)=8
f(2)=-f(1)+6=-2
f(3)=-f(2)+6=8
f(4)=-f(3)+6=-2
f(5)=-f(4)+6=8
Cheers,
Stan H.